This is a probability question regarding punnet squares:

Lets say I do a cross for a sex-linked recessive gene and my results are: 2 girls, 2 boys. Of the boys, only 1 is affected.

The question was: What is the probability the child will be a boy and affected?

The answer was: P(XY) * P(aa) = 1/2*1/2= 1/4

But what if the question was: Of the boys that are born, what is the probability they will be affected? Would the answer be 50% in this case or still 25%?

Similarily, if there were purple:white phenotypes in a 3:1 ratio, and it asked me, of the purple flowers, what is the probability they are homozygous. And 1 out if 3 purple is PP. Then my probability is 1/3? Or is it 1/4(PP)*3/4(All Purple)?

Thank you!
Answer #1

1. Boys that are born - totally 2 boys are borned.
In that 1 will be affected.
So, out of 2 boys, 1 is affected
P = 1/2 = 50%

2. In purple flowers, what is the probability they are homozygous.

Purple flower = PP, Pp, Pp = 3
In that only one is homozygous PP = 1

So, out of 3 only one is homozygous
P = 1/3